A country’s aggregate production function is ? = 50 ×. The depreciation rate is d =...

  1. A country’s aggregate production function is ? = 50 ×8AlWccP7HaNIcAAAAASUVORK5CYII=. The depreciation rate is d = 0.1, population growth rate is n = 0.025, saving rate is s = 0.2.
    1. What is the steady-state output per worker?
    2. What is the steady-state consumption per worker?
    3. What would be the golden steady-state consumption per worker?
Answer #1

Y= AKN ? by N ŽARN y= ASR A-so , d=o. 1 s=o. 2 , n=0.00 Steady state DK=0 Ak=ay - Entd) k = 0 sy - En tok = k 5ok 0.2 0.025toconsumption b) Steady per worker estate (6²) C+B= y c=y-sy c = (1-2yn C= (1-0.2) 4000 (=(0.8) 4004 10% = 32.00] Golda state o